•Given the following code, what is the value of g[5]? The answer is 3.231994 please show me how it is calculated each iteration.

   int i;

   double g[15] = {0.5, 1.5, 2.0, 2.8, 3.2, 0.01};

   double sum = 0;

   for(i = 0; i < 15; i++)

   {

   sum += g[i];

   }

   g[3] = 1.5 + g[0]*g[2];

   g[5] = g[4] + sin( g[4]*g[5]);
Answer #1

From the above code:

g[0] = 0.5

g[1] = 1.5

g[2] = 2.0

g[3] = 2.8

g[4] = 3.2

g[5] = 0.01

In for loop,in each iteration the elements in array g[ ] will add and store in sum.

iteration 1 : i = 0 , g[0] = 0.5 , sum = 0

sum += g[i] means sum = sum + g[i]

sum = sum + g[0] = 0 + 0.5 = 0.5

sum = 0.5

iteration 2 : i = 1 , g[1] = 1.5 , sum = 0.5

sum = sum + g[1] = 0.5 + 1.5 = 2.0

sum = 2.0

iteration 3 : i = 2 , g[2] = 2.0 , sum = 2.0

sum = sum + g[2] = 2.0 + 2.0 = 4.0

sum = 4.0

iteration 4 : i = 3 , g[3] = 2.8 , sum = 4.0

sum = sum + g[3] = 4.0 + 2.8 = 6.8

sum = 6.8

iteration 5 : i = 4 , g[4] = 3.2 , sum = 6.8

sum = sum + g[4] = 6.8 + 3.2 = 10.0

sum = 10.0

iteration 6 : i = 5 , g[5] = 0.01 , sum = 10.0

sum = sum + g[3] = 10.0 + 0.01 = 10.01

sum = 10.01

iteration 7 - 15 :g[6] to g[14] are not initialized so these have value = 0

So, sum will not change,sum = 10.01

g[3] = 1.5 + g[0] * g[2] = 1.5 + (0.5*2.0) = 2.5

g[3] = 2.5

g[5] = g[4] + sin( g[4] * g[5]) -> Assume this as equation 1

Here sin() function return the sine of a number.

g[4] = 3.2 , g[5] = 0.01

sin( g[4] * g[5]) = sin( 3.2 * 0.01) = sin(0.032) = 0.031995 .Substitute this value in  equation 1 ,Then

g[5] = 3.2 + 0.031994 = 3.231994

So g[5] = 3.231994